OpenStudy (anonymous):

linear equation solve. Please help 1/2x+6/5=2/5x My answer was a -12 but it wont check

OpenStudy (anonymous):

6/5 = 2/5x - 1/2x 6/5 = -1/10x -12 = x So you are correct! 1/2(-12)+6/5 = 2/5 (-12) -6 +6/5 = -24/5 -30+6/5=-24/5 -24/5 = -24/5 So it does check!

OpenStudy (anonymous):

thank you, i must have been plugging it in wrong

OpenStudy (anonymous):

when i am multiplying do i have to have a LCD or the same denominator?

OpenStudy (anonymous):

No. When you are multiplying fractions, the denominator does not have to be the same. However, the denominators must be the same for addition.

OpenStudy (anonymous):

how about 17/16x+x/16=6x+1/8+15/16x I got x=-2/93